All Systems Operational
API Operational
Control Panel - https://cloud.glesys.com Operational
Load Balancer Service Operational
VPS Service Operational
Platform: OpenVZ Operational
Platform: VMware Operational
Platform: KVM Operational
Datacenter Falkenberg Operational
Power grid Operational
Core network Operational
Cooling Operational
Edge network Operational
Datacenter Stockholm Operational
Power grid Operational
Network Operational
Cooling Operational
Datacenter Amsterdam Operational
Power Grid Operational
Network Operational
Datacenter Oslo Operational
Power Grid Operational
Network Operational
Datacenter London Operational
Power Grid Operational
Network Operational
Other Operational
Portlane ? Operational
Gothenburg POP Operational
Provider edge Stockholm Operational
Stokab KN7 Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Mar 6, 2021

No incidents reported today.

Mar 5, 2021
Resolved - This incident has been resolved.
Mar 5, 20:35 CET
Monitoring - The packet loss is subsiding, and our network are returning to normal. We will continue to monitor for a while.
Mar 5, 20:24 CET
Investigating - We are seeing packet loss to our services. We are currently investigating the issue.
Mar 5, 19:51 CET
Mar 4, 2021

No incidents reported.

Mar 3, 2021

No incidents reported.

Mar 2, 2021
Resolved - The issue was resolved by the telephone-provider.
Mar 2, 16:51 CET
Investigating - Customers are reporting issues reaching us by phone, we're investigating the issue.
In the meantime, please contact us using support@glesys.se.
This doesn't affect our elevated SLA phone-number, if you have an SLA-pin code you can reach us at 08-525 093 41
Mar 2, 10:47 CET
Mar 1, 2021

No incidents reported.

Feb 28, 2021

No incidents reported.

Feb 27, 2021

No incidents reported.

Feb 26, 2021

No incidents reported.

Feb 25, 2021

No incidents reported.

Feb 24, 2021
Resolved - This incident has been resolved.
Feb 24, 14:07 CET
Monitoring - The packet loss is subsiding, and our network are returning to normal. We will continue to monitor for a while.
Feb 24, 10:30 CET
Identified - The issue has been identified and a fix is being implemented.
Feb 24, 10:21 CET
Update - We are continuing to investigate this issue.
Feb 24, 10:11 CET
Investigating - We are seeing packet loss to our services. We are currently investigating the issue.
Feb 24, 10:07 CET
Feb 23, 2021

No incidents reported.

Feb 22, 2021

No incidents reported.

Feb 21, 2021

No incidents reported.

Feb 20, 2021

No incidents reported.